Suppose two people flip a coin three times. Let X1, X2 denote the number of tails flipped by the first and second person. Find the sampling distribution of the sample mean.
Answer #1

P(X1=0)= 1/2×1/2×1/2 =1/8

P(X1=1)= 3/8

P(X1=2)= 3/8

P(X1=3)= 1/8

Similar distribution for X2

E(x)= xp(x) = 3/8 + 2×3/8 +3×1/8 = 12/8 = 1.5